At takeoff a commercial jet has a 55.0 m/s speed. Its tires have a diameter of 0.700 m. (a) At how many rpm are the tires rotating? Incorrect: Your answer is incorrect. rpm (b) What is the centripetal acceleration at the edge of the tire? Incorrect: Your answer is incorrect. m/s2 (c) With what force must a determined 10-15 kg bacterium cling to the rim? Incorrect: Your answer is incorrect. N (d) Take the ratio of this force to the bacterium's weight. Incorrect: Your answer is incorrect. (force from part (c) / bacterium's weight)
Answer #1

a)

angular velocity

w = v/r = 55 / 0.7 = 78.571 rad/s

w = 78.571* 60 / 2pi = 750.68 rpm

=====

b)centripetal acceleration

ac= r w^2 =0.7* 78.571^2

ac = 4321.43 m/s^2

====

c)

F = ma = 4.321* 10^-12 N

=====

d)

ratio = a /g = 441

=====

comment before rate in case any doubt, will reply for sure. goodluck